The spring oscillates to the other end. The amplitude is 3.5m, k = 1.2, and angular frequency is = 4.3 . Solve the frequency, wavelength, wave speed, period, wave number, and direction.

1
Expert's answer
2021-11-18T15:08:57-0500

wave number: "k=1.2"

angular frequency: "\omega=4.3"


Wavelength:



"\lambda=2\pi/k=5.2\text{ m}."

Frequency:



"f=\omega/2\pi=4.3/2\pi=0.68\text{ Hz}."


Period:



"T=1/f=1.46\text{ s}"



direction: line between ends of spring